OpenStudy (idealist10):

Find a second solution y2 for (2x+1)xy"-2(2x^2-1)y'-4(x+1)y=0; y1=1/x that isn't a constant multiple of the solution y1.

OpenStudy (idealist10):

\[y _{2}=\frac{ u }{ x }\]

OpenStudy (idealist10):

\[y _{2}'=\frac{ u' }{ x }-\frac{ u }{ x^2 }\]

OpenStudy (idealist10):

\[y _{2}''=\frac{ u'' }{ x }-\frac{ 2u' }{ x^2 }+\frac{ 2u }{ x^3 }\]

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

i think its 0

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

Plug in the derivatives: \[\begin{align*} 0&=(2x+1)xy''-2(2x^2-1)y'-4(x+1)y\\\\ &=(2x+1)x\left(\frac{u''}{x}-\frac{2u'}{x^2}+\frac{2u}{x^3}\right)-2(2x^2-1)\left(\frac{u'}{x}-\frac{u}{x^2}\right)-4(x+1)\frac{u}{x}\\\\ &=(2x+1)u''-\frac{2(2x+1)}{x}u'+\frac{2(2x+1)}{x^2}u-\frac{2(2x^2-1)}{x}u'\\ &\quad\quad+\frac{2(2x^2-1)}{x^2}u-4\frac{x+1}{x}u\\\\ &=(2x+1)u''-(4x+4)u' \end{align*}\] Substituting \(t=u'\) gives a linear ODE, \[(2x+1)t'-4(x+1)t=0\]

OpenStudy (idealist10):

Thank you!

OpenStudy (idealist10):

But at the end, I got \[y _{2}=\frac{ u }{ x }=\frac{ C _{1}e ^{x/2}(4x-6)+C _{2} }{ x }\]

OpenStudy (idealist10):

The answer in the book is \[y _{2}=e ^{2x}\].

OpenStudy (anonymous):

Hmm, checking with Mathematica says \(e^{2x}\) should be the solution, but our work thus far gives \((2x+1)e^{2x}\) is... The algebra checks out, so I'm not sure what the problem is.

OpenStudy (idealist10):

So how do we get the right solution e^(2x)?

OpenStudy (anonymous):

I'll look into it and report back when I find out :)

OpenStudy (idealist10):

Okay.

OpenStudy (anonymous):

It doesn't look like any of my work is wrong, so I'm thinking you made a mistake somewhere when you solved for \(u\). Let's return to the linear ODE in terms of \(t=u'\): \[(2x+1)t-4(x+1)t=0~~\iff~~t'-4\frac{x+1}{2x+1}t=0\] Separating the variables gives \[\frac{dt}{t}=4\frac{x+1}{2x+1}\,dx\] and integrating gives \[\ln|t|=(2x+1)+\ln|2x+1|+C_1~~\implies~~t=C_1(2x+1)e^{2x+1}\] Integrating again (to solve for \(u\)) gives \[u=\int C_1(2x+1)e^{2x+1}\,dx=\frac{C_1}{2}(2x+1)e^{2x+1}-\frac{C_1}{2}e^{2x+1}+C_2\] which reduces to \[u=C_1xe^{2x+1}+C_2=C_1xe^{2x}+C_2\] Dividing by \(x\) to solve for \(y_2\) gives what we want: \[y=\frac{u}{x}~~\implies~~y_2=e^{2x}\]

OpenStudy (idealist10):

Thank you! I finally got this problem!